OpenStudy (anonymous):

Solve each equation by completing the square. a^2+16-66=-7

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

ok,add 7 to both sides first :)

OpenStudy (anonymous):

done

OpenStudy (anonymous):

ok you have \(\sf \large a ^{2} + 16 - 66 + 7 = 0\)

OpenStudy (anonymous):

can you say this? 16 - 66 + 7 = ?

OpenStudy (anonymous):

yes I do.

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

16 - 66 + 7 = -43

OpenStudy (anonymous):

so it's \( \sf \large a ^{2} - 43 = 0\)

OpenStudy (anonymous):

you can solve it by using the formula,or you can solve it by using the laws of mathematics :)

OpenStudy (anonymous):

the standard form of these equations are: \(\sf \Large Aa ^{2} + Ba + C = 0\)

OpenStudy (anonymous):

and the formula is: \(\sf \LARGE a _{1,2} = \frac{ -b \ \pm \ \sqrt{b ^{2}-4ac}}{ 2a }\)

OpenStudy (anonymous):

can you do it yourself?

OpenStudy (anonymous):

so just use the quadratic formula to solve?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

okay thankyou. I think I can take it from there. :)

OpenStudy (anonymous):

and do you want to know how to solve it by the laws of mathematics? besides that,do you want the proof of this formula?

OpenStudy (anonymous):

what does that mean? sorry I'm not good at math...or english... to be honest.

OpenStudy (anonymous):

oh no,no problem :) i'm not good at english too :) it mean: \(\ \sf \large Every \; equations \; can \; be \; solved \; by \; that \; formula \; or \; by \; the \; laws \; of \; mathematics \) what does "by the laws of mathematics" mean? i'll tell you :)

OpenStudy (anonymous):

\( \sf a ^{2} - 43 = 0\) so can we say \( \sf a ^{2} = 43\) ?

OpenStudy (anonymous):

yes we can.

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

nice :) please solve this,then you can find you answer: \(\sf a ^{2} = 4\) solve for \( \sf a \)

OpenStudy (anonymous):

a=2. right?

OpenStudy (anonymous):

just 2?

OpenStudy (anonymous):

\(\sf 2 ^{2} = 4\) it's right

OpenStudy (anonymous):

but \(\sf(-2)^{2} = 4\) too :)

OpenStudy (wolf1728):

A quadratic function can be solved by • factoring • the quadratic formula OR • completing the square Knowing 1 of these methods will allow you to solve all quadratic equations but it seems algebra teachers want to show all 3 ways.

OpenStudy (anonymous):

how can you solve\(\sf a^{2} = 4 \) by factoring ?:D

OpenStudy (wolf1728):

a² - 4 = 0

OpenStudy (anonymous):

it's factoring ???

OpenStudy (wolf1728):

You would be factoring the difference of 2 squares

OpenStudy (anonymous):

yes you can say this,but i have another meaning of "Factoring" ... and now @Mhm120 , got it? so \(\sf (-2)^{2} = 4\) too

OpenStudy (wolf1728):

Now that I think of it, trying to factor 3x² -17 +23 = 0 would be extremely difficult.

OpenStudy (anonymous):

@wolf1728 , we're just confusing him . wait please.

OpenStudy (wolf1728):

Sure as a matter of fact, I think I'll just leave this topic.

OpenStudy (anonymous):

no,no your quite right :) but he know knows the formula,and i wanted to tell him how to solve that by another way, and you're quite right too.just i said show your work after me :)